OpenStudy (anonymous):

What is the area of a regular polygon when one side equals 6m?

OpenStudy (anonymous):

well...

OpenStudy (anonymous):

???

OpenStudy (anonymous):

it depend on the number of sides of the polygon

OpenStudy (anonymous):

*depends

OpenStudy (anonymous):

Oh I think its a Pentagon

OpenStudy (anonymous):

Would it be 61.94?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\sqrt{5(5+2\sqrt{5})} a ^{2}\]

OpenStudy (anonymous):

that's the formula